OpenStudy (anonymous):

MEDALS Gene has a gasoline budget of $275 per month. He uses an average of $5 of gasoline each day he drives. Which of the following equations represents how much money is left in his gasoline budget after x days of driving? A. y = 275 - 5x B. y = 275 + 5x C. y = 275x - 5 D. y = 5x - 275

OpenStudy (anonymous):

y = 275 - 5x is the answer i suppose

OpenStudy (anonymous):

suppose he drove 5 day & average of one day is $5 so 5 * 5 = 25 275 - 25 = 250 $ is left

OpenStudy (anonymous):

Thank you :D can I get a chocolate chip cookie ^_^